### v1

# OccFiner: Offboard Occupancy Refinement with Hybrid Propagation

Abstract
Vision-based occupancy prediction, also known as 3D Semantic Scene Completion (SSC), presents a significant challenge in computer vision. Previous methods, confined to onboard processing, struggle with simultaneous geometric and semantic estimation, continuity across varying viewpoints, and single-view occlusion. Our paper introduces OccFiner, a novel offboard framework designed to enhance the accuracy of vision-based occupancy predictions. OccFiner operates in two hybrid phases: 1) a multi-to-multi local propagation network that implicitly aligns and processes multiple local frames for correcting onboard model errors and consistently enhancing occupancy accuracy across all distances. 2) the region-centric global propagation, focuses on refining labels using explicit multi-view geometry and integrating sensor bias, especially to increase the accuracy of distant occupied voxels. Extensive experiments demonstrate that OccFiner improves both geometric and semantic accuracy across various types of coarse occupancy, setting a new state-of-the-art performance on the SemanticKITTI dataset. Notably, OccFiner elevates vision-based SSC models to a level even surpassing that of LiDAR-based onboard SSC models.
Keywords:
Semantic Scene Completion 3D Occupancy Prediction Offboard Perception
1 Introduction
Figure 1: Current onboard methods generate unreliable occupancy predictions that are inconsistent across different viewpoints. In contrast, our offboard framework constructs a unified and multi-view consistent occupancy map with higher accuracy.
Vision-based Occupancy Prediction, often referred to as Semantic Scene Completion (SSC), is a process aimed at accurately reconstructing the geometry and semantics of a surrounding 3D scene from the image captures [6]. Compared to LiDAR solutions, camera-based systems offer the benefits of being lightweight, cost-effective, and ease of deployment and maintenance. Most SSC efforts are made in onboard settings [52, 6, 27], i.e., data calculation and feedback take place on the mounted vehicle. However, integrating data into offboard systems creates opportunities for identifying and rectifying potential deficiencies in models. Consequently, we propose offboard SSC, a new setup that entails additional computation and refinement of on-vehicle predictions at a data center. It can play a pivotal role in auto-labeling by leveraging aggregated and diverse data from multiple sources, significantly reducing manual annotation efforts and improving labeling accuracy for large-scale 3D scene understanding [77, 25, 24].
Despite various research efforts [6, 27, 18], vision-based SSC has fallen short in accuracy compared to LiDAR solutions. There are two limitations: 1) The depth measurement of monocular cameras lacks the precision of LiDAR systems, requiring onboard vision-based SSC methods to struggle with both geometry and semantics. 2) Onboard models are limited to processing a restricted number of local frames, impacting accuracy and multi-view continuity in both vision- and LiDAR-based SSC. This issue manifests as unreliable estimations and discontinuities across viewpoints in geometry and semantics, as illustrated in Fig. 1. We argue that high-fidelity SSC maps should remain consistent across viewpoints to ensure safety and reliability, especially since static rigid objects should not exhibit significant changes in classes. While offboard systems offer refined evaluation of onboard model performance, essential for the data closed-loop in autonomous driving [77, 25, 24], their potential in SSC remains under-explored.
Given such limitations of inferior performance and onboard SSC setting, we introduce OccFiner, the first offboard occupancy refinement framework. The objective is to significantly enhance the reliability of vision-based SSC. By doing so, the constructed 3D SSC map can be reliably used repeatedly after a one-time low-cost acquisition with only cameras. Specifically, we consider two primary error sources in addressing the inaccuracies of vision-based SSC: the prediction bias error introduced by the onboard model, and the measurement bias from the camera capture process. OccFiner decouples these two heterogeneous error sources and uses a two-stage hybrid process to process them separately: (1) Multi-to-Multi Local Propagation. In the initial stage, OccFiner addresses prediction bias errors from the onboard model. We implement an error compensation model that processes inputs from off-the-shelf SSC models [6, 27, 52]. This model uniquely encodes relative spatial coordinates within a local window, facilitating the propagation of geometric and semantic cues by a newly proposed DualFlow4D transformer. This approach effectively enhances the accuracy of occupancy predictions at varying distances. (2) Region-centric Global Propagation. The second stage focuses on measurement errors from camera captures. Here, OccFiner employs explicit multi-view geometry to register and aggregate data across the entire scene, incorporating senor bias weighting in the process. Compared with the simple accumulation of multi-frame averages, our region-centric voting strategy yields a more accurate SSC map. By integrating local implicit error compensation and non-local explicit information propagation from various viewpoints, OccFiner effectively addresses onboard bias, multi-view consistency, and single-view occlusions, therefore naturally enhancing the reliability and robustness of scene understanding.
To the best of our knowledge, we are the first to introduce offboard refinement for SSC. Our method, termed OccFiner, enables seamless integration across various SSC models and datasets as a plug-and-play solution. For 3D semantic scene completion on SemanticKITTI dataset [2], OccFiner has demonstrated remarkable performance, surpassing VoxFormer by $3.67\%$ in mIoU, which makes a $26.99\%$ relative improvement and ranks first on the test leaderboard among all monocular methods. Surprisingly, OccRefiner also works well for LiDAR-based SSC [65] with minor changes, leading to a new state-of-the-art of $37.82\%$ in mIoU, achieving a $9.5\%$ improvement compared with the best benchmark result.
To summarize, we deliver the following contributions:
  • We are the first to explore the problem of learning to generate high-quality vision-oriented SSC maps offboard.
  • We propose OccFiner, an effective offboard occupancy refinement framework with hybrid propagation to build multi-view consistent SSC maps.
  • OccFiner sets new state-of-the-art performances in both camera-based and LiDAR-based SSC. For the first time, the accuracy of the camera-based SSC method surpasses that of the classical onboard LiDAR-based method [55].
2 Related Work
2.1 3D Semantic Scene Completion
Semantic Scene Completion (SSC), also known as Semantic Occupancy Prediction, aims at concurrently estimating the geometry and semantics of a surrounding scene. The first effort SSCNet [55] lays the foundation for this field by defining the SSC task and introducing a method for estimating the structure and semantics of indoor scenes from a singular depth image input. Recognizing the critical role of SSC-based perception in the realm of autonomous driving, researchers have increasingly turned their attention to its applications in outdoor scenes after the release of the large-scale outdoor benchmark SemanticKITTI [2]. The predominant works in this area can be broadly categorized based on the input modality: LiDAR-based SSC [52, 51, 70, 8, 73, 85, 65, 69] and vision-based SSC [53, 36, 38, 63, 34, 64, 35, 33, 20, 75, 29, 6, 82, 16]. Additionally, recent research expands into several novel areas within SSC including the world model [83], 3D object detection [19, 59, 49, 68, 15, 84, 41], multimodal parsing [22, 74, 79, 45, 39, 42, 66, 31, 43], self-supervised prediction [14, 78, 40, 17, 4], open-vocabulary recognition [60, 57], dense top-view understanding [48, 12, 32], V2X collaboration [28, 81, 54], as well as various benchmarks [62, 58, 26]. However, the majority of previous studies have concentrated on methodological design within onboard settings, where observations across different viewpoints often exhibit notable discontinuities. In contrast, OccFiner shifts focus to the offboard setting, aiming to rectify errors inherent in onboard models and to aggregate long-term geometric and semantic cues. Our approach contributes to data closure in autonomous driving, addressing a crucial aspect that has been largely overlooked in previous SSC research.
2.2 Offboard 3D Perception
Perception in autonomous driving scenarios necessitates extensive data annotation and training [11, 5, 56]. Since LiDAR can provide accurate range information in 3D space, researchers explore its sequence learning in semantic segmentation [61, 1], object detection [44, 47] and tracking [46] to obtain better performance. The sequential characteristics of LiDAR scans are further studied in offboard 3D object detection [50, 71, 10, 37] to reduce annotation costs. 3DAL [50] formulates the problem of offboard 3D object detection and proposes an object-centric auto-labeling detector. CTRL [10] adheres to the principle of track-centric and elevates the accuracy of auto-labeling, surpassing manual annotation in some scenarios. The aforementioned methods are all designed for LiDAR point cloud, attributed to the ease of accumulating 3D spatial information compared to images. MV-Map [67] provides a region-centric fusion solution for high-definition maps with camera data as inputs while it only processes limited traffic elements on the bird’s eye view plane. To address the above challenges, We design the first hybrid propagation pipeline (OccFiner) for offboard semantic scene completion (SSC). OccFiner can process the SSC predictions from either cameras or LiDAR and provide high-quality occupancy information for autonomous vehicles, which holds significant importance in auto-labeling and data closure in pure visual perception.
3 Approach
3.1 Overview
The OccFiner framework, designed for offboard Semantic Scene Completion (SSC), operates in two distinct stages. The first stage focuses on compensating for errors in onboard model predictions, employing a multi-to-multi local propagation network to process SSC model outputs across a series of frames, utilizing both close-range and distant frames as references. This stage effectively merges relative spatial coordinates with semantic features, employing transformer-based methods for spatio-temporal integration. The second stage of OccFiner emphasizes global semantic and geometric aggregation. It transforms refined voxel labels into semantic point clouds and applies relative poses over extended periods for precise coordinate adjustments, incorporating sensor measurement characteristics for voxel voting. This dual-stage strategy effectively combines local implicit feature propagation with global explicit semantic aggregation, forming a robust hybrid system for offboard SSC refinement.
Figure 2: Overview of the proposed multi-to-multi local propagation network. It accepts multiple onboard predictions and relative coordinates as input. This network adeptly executes error compensation and facilitates the implicit local propagation, improving SSC quality across various distances.
3.2 Multi-to-Multi Local Propagation Network
Problem formulation: Let $X_{1}^{T}:=\{X_{1},X_{2},...,X_{T}\}$ be a onboard SSC inference sequence of height $H$, width $W$, depth $Z$, and frames length $T$. $C_{1}^{T}:=\{C_{1},C_{2},...,C_{T}\}$ denotes the corresponding frame-wise relative coordinates, which can be calculated by:
$\mathbf{c}_{i}=(T_{li}^{cam})^{\intercal}(T_{cam_{t}}^{world})^{\intercal}T_{cam_{i}}^{world}T_{li}^{cam}\mathbf{x}_{i},$ (1)
where each vertex $\mathbf{x}_{i}\in X_{i}$ is projectively associated to a relative vertex $\mathbf{c}_{i}\in C_{i}$. For each relative coordinate $C_{i}$, it represents the mapping relationship between the three-dimensional coordinates of the current frame $X_{i}$ and the coordinates of the pivot frame $X_{t}$. We formulate the occupancy refinement as a task that takes the onboard ($X_{1}^{T},C_{1}^{T}$) pairs as input and reconstruct the SSC labels $Y_{1}^{T}=\{Y_{1},Y_{2},...,Y_{T}\}$. Specifically, we propose to learn a mapping function from onboard inferences $X_{1}^{T}$ to the output $\hat{Y}_{1}^{T}:=\{\hat{Y}_{1},\hat{Y}_{2},...,\hat{Y}_{T}\}$, such that the conditional distribution of the real data $p(Y_{1}^{T}|X_{1}^{T})$ can be approximated by the one of generated data $p(\hat{Y}_{1}^{T}|X_{1}^{T})$.
We conceptualize the task as a “multi-to-multi” prediction problem, focusing on the simultaneous refinement of all input SSC frames in a single feed-forward process. The intuition is that regions obscured in the current viewpoint are likely to become visible in frames captured from a distance, particularly in scenarios involving large occlusions or when the vehicle is moving at a slower pace. In the context of multi-frame offboard SSC, it is more effective to address gaps in a target frame by leveraging content from the entire scene sequence, incorporating information from both neighboring and distant frames as conditional inputs. This approach relies on the Markov assumption [13] for simplification, which allows us to express the refinement process as:
$p(\hat{Y}_{1}^{T}|X_{1}^{T})=\prod_{t=1}^{T}p(\hat{Y}_{t-n}^{t+n}|X_{t-n}^{t+n},X_{t-d,s}^{t+d}),$ (2)
where $X_{t-n}^{t+n}$ represents a short SSC clip of neighboring frames centered around time $t$ with a temporal radius $n$. $X_{t-d,s}^{t+d}$ denotes distant frames that are randomly sampled from a distant scene range $d$ at a rate of $s$. This selection of distant frames typically encompasses most of the key moments in possible viewpoints, effectively conveying long-term geometric and semantic cues of the scene.
Within this framework, offboard SSC models are tasked with not only maintaining temporal consistency in neighboring frames but also ensuring that the refined frames are coherent with the broader narrative of the scene sequence. This approach aims to enhance scene understanding by integrating both local and global temporal perspectives in the refinement process.
Network design: The overview of the proposed multi-to-multi local propagation network is shown in Fig. 2. As indicated in Eq. 2, OccFiner takes both neighboring local SSC frames $X_{t-n}^{t+n}$, distant reference SSC frames $X_{t-d,s}^{t+d}$, and relative coordinate $\{C_{t-n}^{t+n},C_{t-d,s}^{t+d}\}$ as conditions, to propagate across all input frames simultaneously. Specifically, OccFiner is comprised of four integral components, including a frame-level BEV feature encoder, a parallel BEV positional encoder, a series of spatial-temporal transformers, and a frame-level segmentation head. The frame-level BEV feature encoder is architecturally configured with multiple 2D convolutions along the horizontal and vertical dimensions (X, Y), thus converting the height dimension (Z) into a feature dimension. Mirroring this, the frame-level BEV positional encoder shares the same architectural blueprint with the feature encoder, such a design choice aids in better aligning the positional and feature information, as both are processed through a similar computational pathway. At the heart of OccFiner lie the DualFlow4D transformers, tasked with learning and applying joint spatial-temporal transformations. These transformers are specifically tailored to address uncertainties and occlusions within the deep encoding space.
Figure 3: Our proposed DualFlow4D transformer block. It engages spatiotemporal propagation within BEV space and vanilla attention to pillar tokens. This dual approach enables effective matching and flow of semantic and geometric cues for comprehensive scene understanding.
DualFlow4D transformer: To propagate high-fidelity features in each SSC frame, multi-layer DualFlow4D transformers are designed to search coherent contents from all the encoded features. Specifically, we propose to first search by a multi-head soft-patch-based focal attention module along BEV and temporal dimensions. We use $f_{1}^{T}=\{f_{1},f_{2},...,f_{T}\}$, where $f_{i}\in\mathbb{R}^{h\times w\times c}$ to denote the features encoded from the frame-level BEV encoder. Concurrently, $p_{l}\in\mathbb{R}^{T_{l}\times h\times w\times c}$ and $p_{r}\in\mathbb{R}^{T_{r}\times h\times w\times c}$ encapsulate the corresponding encoded local and reference relative coordinates. We first use soft split to embed them into overlapped patches $P\in\mathbb{R}^{(T_{l}+T_{r})\times W_{h}\times W_{w}\times c_{e}}$:
$P={\rm SS}(({f}_{l}\oplus f_{r})+({p}_{l}\oplus p_{r})),$ (3)
where ${\rm SS}(\cdot)$ denotes the soft split operation, which softly splits each BEV feature into overlapped patches of size $W_{h}\times W_{w}\times T$ with stride $k$, and flattened to a one-dimensional token, which is similar to the image splitting strategy in T2T-ViT [76]. $T_{l}$ and $T_{r}$ are the time dimension of local frames and reference frames, and $W_{h}$ and $W_{w}$ are the spatial dimension of embedded tokens. $\oplus$ denotes the feature concatenation. The overlapped position aggregates a piece of information from different tokens, contributing to smoother semantic patch boundaries and enlarging its receptive field by fusing cues from neighboring patches.
Instead of the vanilla vision transformer [9], we use focal transformer [72] to search from both local and non-local neighbors to propagate high-fidelity BEV features. Concretely, $P$ is linearly projected to queries $Q$, keys $K$, and values $V$ for computing the 3D focal attention:
$Q,K,V=\mathcal{P}_{qkv}({\rm LN}(P)),$ (4)
$Z_{bev}={\rm MHFA}(Q,K,V)+P,$
where ${\rm LN}$ and ${\rm MHFA}$ denote the layer normalization and multi-head focal attention, respectively, $\mathcal{P}_{qkv}$ is the linear projection layer. We omit the time dimension for simplicity. The previous propagation is limited to the temporal BEV space. To ensure effective propagation in the height direction, we have additionally introduced attention along the z-axis, specifically focusing on the ‘pillar’, which refers to a voxel grid oriented along the height:
$Z_{z}:={\rm Attn}(Q_{z},K_{z},V_{z})={\rm Softmax}(\frac{Q_{z}K_{z}^{\top}V_{z}}{\sqrt{d}}),$ (5)
where $Q_{z},K_{z},V_{z}\in\mathbb{R}^{W_{z}\times T\times c_{e}}$ are respectively reshaped queries, keys, and values. Finally, the output of the entire propagated feature will be gathered:
$Z_{4d}=\mathcal{P}_{z}(Z_{z})+\mathcal{P}_{bev}(Z_{bev}),$ (6)
where $\mathcal{P}_{z}$ and $\mathcal{P}_{bev}$ are linear projections. Note that the above formulas omit the head dimension for simplicity. Such a 4D dual-branch design promotes learning coherent spatial-temporal transformations by preserving both BEV information flow and the height dimension information flow along the pillar-shape voxels.
Figure 4: Overview of the proposed region-centric global propagation. Given an arbitrary number of input frames, we leverage multi-view geometry and consider sensor bias to augment long-term SSC consistency and accuracy.
3.3 Region-centric Global Propagation
In the initial stage, the OccFiner framework engages in multi-to-multi local propagation for offboard SSC (Semantic Scene Completion) refinement. This stage, however, is limited by the input multi-frame window size, leading to a necessity for global propagation across the entire scene sequence. Notably, the local propagation process uniformly enhances SSC accuracy regardless of the various distances, a key aspect of global propagation’s effectiveness. Additionally, OccFiner utilizes relative coordinates in the initial stage, ensuring the implicit alignment of multiple frames and maintaining the contextual integrity of each frame’s semantic content. This approach, which prevents truncation, is pivotal for learning-based error compensation. The subsequent stage involves geometric explicit registration voting, covering the entire scene sequence, and forming a cohesive and hybrid strategy within the OccFiner framework. In the global propagation stage, OccFiner focuses on addressing the measurement limitations of camera and LiDAR sensors and aggregates the independent SSC predictions into a multi-view consistent prediction for each region. OccFiner adeptly integrates sensor-aware weight biases with vanilla registration methods, aligning the refinement process with the unique attributes of each sensor type. Next, we delve into the details of vanilla registration and enhanced sensor-aware registration.
Vanilla Registration: As illustrated in Fig. 4, it aggregates independent SSC predictions from $12n+$ 1 frames into a multi-view consistent prediction for each region, where $n$ is the temporal window radius. This involves the aggregation of per-frame semantics and geometry $\{\hat{Y}_{i}\}_{i=1}^{2n+1}$ into a refined SSC map. For a given target voxel index $\mathbf{y}$ within a 3D region of target $\hat{Y}_{tg}$, the process starts with devoxelizing the current SSC voxel $\hat{Y}_{cur}$ to derive the semantic point cloud $\mathbf{P}_{cur}$ in the current LiDAR coordinate system:
$\mathbf{P}_{cur}(x,y,z,c)=\left\{\left(o_{x}+i\cdot\text{dv}_{x},o_{y}+j\cdot\text{dv}_{y},o_{z}+k\cdot\text{dv}_{z},c\right)\,|\,\hat{Y}_{ijk}=c\right\},$ (7)
where the voxel origin is $(o_{x},o_{y},o_{z})$. dv denotes the voxel size and $c$ represents the semantic classes. This point cloud is then transformed relative to the target LiDAR coordinate system with poses $\{{T}_{cur}^{tg}\}_{tg-n}^{tg+n}$:
$\mathbf{P}_{cur}^{tg}={T}_{cur}^{tg}\mathbf{P}_{cur},$ (8)
followed by voxelization and addition in the voting voxel. The final step involves voting for the final semantic voxel output $\overline{Y}_{final}$:
[…]
where the $\mathrm{argmax}$ function selects the class $c$ with the highest aggregated value in the voting voxel, determining the final semantic voxel.
Incorporating with Sensor Bias: While vanilla registration averages SSC across multiple frames, it overlooks the differential weighting of different locations. To address this, sensor-aware weighting is introduced, tailored to the distinct characteristics of cameras and LiDARs. This approach is informed by the measurement characteristics of measurement devices like cameras and LiDARs. Camera systems are constrained by their field of view, with points within the viewing frustum yielding higher accuracy compared to those outside. Additionally, the occlusion of distant points by nearer ones necessitates lower voting weights for the former. Conversely, LiDAR systems exhibit increased sparsity and depth measurement errors at longer distances. Therefore, for cameras, points within the field of view and at closer ranges are given higher voting weights:
$F_{cam}(x_{c},y_{c},z_{c})=\left(|\theta_{xc}|\leq\frac{fov_{h}}{2}\right)\land\left(|\theta_{yc}|\leq\frac{fov_{w}}{2}\right)\land(z_{c}>0),$ (10)
$W_{cam}(\mathbf{x})=\begin{cases}w_{\text{high}},&\text{if $(F_{cam}\land B)(\mathbf{x})>0$}\\ w_{\text{med}},&\text{if $[F_{cam}(\mathbf{x})>0]$ $\land$ $[B(\mathbf{x})<0]$}\\ w_{\text{low}},&\text{otherwise}\end{cases}$ (11)
where $(x_{c},y_{c},z_{c})$ are the voxel indices in the camera coordinate system, $fov$ denotes the field of view of the camera, $F_{cam}$ represents the voxels within the frustum, and $B$ is the bounding box voxel with values being True inside the box and False outside. $\mathbf{x}$ is the vertex and $W_{cam}$ is the final camera voting weight.
While for LiDARs, the voting weight linearly attenuates with increasing distance from the LiDAR center:
$W_{li}(r)=w_{\text{max}}-(w_{\text{max}}-w_{\text{min}})\times\frac{r}{R},$ (12)
where $r$ denotes the radial distance from the origin in the LiDAR coordinate system, $R$ represents the maximum range, and $w_{\text{max}}$ and $w_{\text{min}}$ correspond to the maximum and minimum weights, respectively. Consequently, the OccFiner framework adopts a hybrid dual-stage strategy: the first stage addresses onboard model errors, enhancing prediction accuracy regardless of the range, and the second stage adjusts for the physical limitations of cameras and LiDARs, further improving the offboard accuracy.
3.4 Supervision
Our multi-to-multi local propagation network is trained with the labels from the training set frame-by-frame, while the region-centric global propagation is a learning-free process. The SSC loss $\mathcal{L}_{ssc}$ for network training consists of two parts as below:
$\mathcal{L}_{ssc}=\mathcal{L}_{ce}+\mathcal{L}_{lovasz},$ (13)
where $\mathcal{L}_{ce}$ is the cross-entropy loss and $\mathcal{L}_{lovasz}$ is the Lovász loss [3].
4 Experiments
4.1 Datasets and Implementation Details
SemanticKITTI and SSCBench-KITTI360: The SemanticKITTI dataset [2] is a pivotal SSC benchmark featuring $22$ outdoor driving scenarios, segmented into train, validation, and test sets with a $10/1/11$ split. The focus is on a specific volume around the vehicle, mapped in $256\times 256\times 32$ voxel grids, where each voxel measures $0.2m^{3}$. SSCBench-KITTI360 [26], derived from KITTI-360 [30], spans $73.7km$ with extensive image and laser scan data, providing diverse geographical coverage and extended sequences. This dataset, sampled to reduce redundancy, offers around $13k$ frames, serving as a vital resource for SSC research.
Training and Optimization: Models are trained for $10$ epochs on the SemanticKITTI and $15$ on SSCBench-KITTI360, using the Adam optimizer [21] with a learning rate of $0.001$. The rate decay is $0.98^{epoch}$. Training is performed on $8$ RTX 3090 GPUs and includes x-y flipping augmentation in 3D volume space.
4.2 Comparison with State-of-the-Art
Table 1: Camera-based semantic scene completion results on the SemanticKITTI validation set [2]. * represents these methods are adapted for the RGB inputs, which are implemented and reported in MonoScene [6]. normal-$\dagger$ represents the reproduced result from [18].
Method
$\blacksquare$ car (3.92%)
$\blacksquare$ bicycle (0.03%)
$\blacksquare$ motorcycle (0.03%)
$\blacksquare$ truck (0.16%)
$\blacksquare$ other-vehicle (0.20%)
$\blacksquare$ person (0.07%)
$\blacksquare$ bicyclist (0.07%)
$\blacksquare$ motorcyclist (0.05%)
$\blacksquare$ road (15.30%)
$\blacksquare$ parking (1.12%)
$\blacksquare$ sidewalk (11.13%)
$\blacksquare$ other-ground (0.56%)
$\blacksquare$ building (14.1%)
$\blacksquare$ fence (3.90%)
$\blacksquare$ vegetation (39.3%)
$\blacksquare$ trunk (0.51%)
$\blacksquare$ terrain (9.17%)
$\blacksquare$ pole (0.29%)
$\blacksquare$ traffic-sign (0.08%)
IoU mIoU
LMSCNet* [52] 18.33 0.00 0.00 0.00 0.00 0.00 0.00 0.00 40.68 4.38 18.22 0.00 10.31 1.21 13.66 0.02 20.54 0.00 0.00 28.61 6.70
3DSketch* [7] 18.59 0.00 0.00 0.00 0.00 0.00 0.00 0.00 41.32 0.00 21.63 0.00 14.81 0.73 19.09 0.00 26.40 0.00 0.00 33.30 7.50
AICNet* [23] 14.71 0.00 0.00 4.53 0.00 0.00 0.00 0.00 43.55 11.97 20.55 0.07 12.94 2.52 15.37 2.90 28.71 0.06 0.00 29.59 8.31
JS3C-Net* [70] 24.65 0.00 0.00 4.41 6.15 0.67 0.27 0.00 50.49 11.94 23.74 0.07 15.03 3.94 18.11 4.33 26.86 3.77 1.45 38.98 10.31
OccFormer [82] 25.09 0.81 1.19 25.53 8.52 2.78 2.82 0.00 58.85 19.61 26.88 0.31 14.40 5.61 19.63 3.93 32.62 4.26 2.86 36.50 13.46
[…]
TPVFormer [18] 23.81 0.36 0.05 8.08 4.35 0.51 0.89 0.00 56.50 20.60 25.87 0.85 13.88 5.94 16.92 2.26 30.38 3.14 1.52 35.61 11.36
VoxFormer [27] 27.01 1.05 0.47 9.90 4.64 1.51 0.85 0.00 54.67 18.55 27.35 0.41 19.65 8.52 26.18 6.58 32.39 8.69 4.71 44.16 13.33
Average Fusion (Vox) 30.35 0.88 0.41 12.92 3.99 1.60 0.86 0.00 57.60 18.53 29.44 0.28 21.83 10.50 29.24 7.44 34.61 9.79 5.54 46.03 14.52
OccFiner (Mono) 29.12 0.41 1.41 21.19 7.85 3.97 2.28 0.00 56.62 19.64 27.62 1.48 15.07 7.18 20.89 3.89 31.79 3.92 2.39 37.42 13.45
w.r.t. MonoScene +5.86 -0.20 +0.96 +14.21 +6.37 +2.11 +1.08 - +0.10 +5.37 +0.90 +1.02 +0.98 +1.34 +3.00 +1.08 +2.15 -0.22 +0.14 +0.56 +2.37
OccFiner (TPV) 29.19 0.69 0.00 22.73 7.05 0.95 1.40 0.00 57.15 22.78 28.35 7.93 15.10 6.87 19.88 3.09 31.02 3.70 1.92 36.54 13.30
w.r.t. TPVFormer +5.38 +0.33 -0.05 +14.65 +2.70 +0.44 +0.51 - +0.65 +2.18 +2.48 +7.08 +1.22 +0.93 +2.96 +0.83 +0.64 +0.56 +0.40 +0.93 +1.94
OccFiner (Vox) 36.78 1.73 0.29 32.96 5.67 4.06 1.12 0.00 66.34 27.05 35.06 0.35 25.46 11.79 32.82 9.87 40.56 8.02 3.58 47.86 18.09
w.r.t. VoxFormer +9.77 +0.68 -0.18 +22.05 +1.03 +2.55 +0.27 - +11.67 +8.50 +7.71 -0.06 +5.81 +3.27 +6.64 +3.29 +8.17 -0.67 -1.13 +3.70 +4.76
As our work presents the first offboard SSC generation, there are no existing competitors to parallel OccFiner’s performance. Additionally, our OccFiner can utilize any off-the-shelf onboard model as its intermediate component. As shown in Tab. 1, we use the camera-based SSC algorithm [6, 18, 27] as the onboard component. On the SemanticKITTI validation set [2], OccFiner demonstrates substantial improvements in mIoU accuracy for various algorithms like MonoScene [6], TPVFormer [18], and VoxFormer [27]. The mIoU accuracy is relatively improved by $21.39\%$, $17.08\%$, and $35.71\%$, respectively. Notably, OccFiner (Vox) achieves the new state-of-the-art camera SSC accuracy of $18.09\%$ mIoU. To ensure a meaningful and fair comparison, we also establish an offboard baseline algorithm, “Average Fusion”. It performs region-centric aggregation without learning refinement or considering sensor measurement biases. Despite this, compared to VoxFormer + Average Fusion, our OccFiner (Vox) can still improve SSC quality by a large margin of over $24.59\%$ in mIoU under the long-range setting.
Figure 5: Qualitative comparisons. Our offboard solution effectively fixes critical errors in onboard component [27], such as large areas of missing road and vehicles. Moreover, OccFiner elevates the performance of pure visual solutions beyond the classic LiDAR-based onboard method SSCNet-full [55].
Qualitative comparisons. As shown in Fig. 5, our offboard solution, OccFiner, effectively corrects major onboard errors, notably in resolving extensive missing road and vehicle voxels. Furthermore, OccFiner enhances the effectiveness of purely visual approaches, surpassing the traditional LiDAR-based onboard method SSCNet-full [55] in performance, indicating its potential to advance the field of autonomous navigation with pure visual solution.
Table 2: Quantitative comparison against the state-of-the-art LiDAR-based SSC methods on the SemanticKITTI validation set. normal-$\dagger$ results from the official code release.
Methods Modality IoU (%) mIoU (%) Rel.
12.8m 25.6m 51.2m 12.8m 25.6m 51.2m
VoxFormer [27] Camera 65.38 57.70 44.16 22.13 18.58 13.33 -
OccFiner (VoxFormer) Camera 65.42 58.50 47.86 23.43 21.29 18.09 $\uparrow$24.59%
SSCNet [55] LiDAR 50.85 52.54 44.90 17.59 17.39 14.11 -
OccFiner (SSCNet) LiDAR 69.64 67.65 55.92 23.90 23.26 19.28 $\uparrow$36.64%
SSCNet-full [55] LiDAR 64.37 61.02 50.22 20.02 19.68 16.35 -
OccFiner (SSCNet-full) LiDAR 65.96 64.55 54.45 23.41 23.31 19.77 $\uparrow$20.92%
LMSCNet [52] LiDAR 74.88 69.45 55.22 22.37 21.50 17.19 -
OccFiner (LMSC) LiDAR 75.18 69.72 57.25 24.01 23.29 19.88 $\uparrow$15.65%
SCPNetnormal-$\dagger$ [65] LiDAR 73.88 64.09 49.06 48.62 44.62 35.06 -
OccFiner (SCPNet) LiDAR 76.44 67.33 57.19 48.70 45.90 40.26 $\uparrow$14.83%
Our exploration extends to validating OccFiner’s versatility across different modalities in offboard SSC generation. As shown in Tab. 2, for LiDAR-based SSC models like SSCNet [55], SSCNet-full [55], LMSCNet [52], and SCPNet [65], OccFiner consistently attains relative mIoU improvements of $36.64\%$, $20.92\%$, $15.65\%$, and $14.83\%$, respectively. This demonstrates its modality-agnostic capability. We observe that the enhancements offered by OccFiner are consistent across varying distances. This improvement is not only confined to long distances and low-accuracy scenarios but also extends to challenging high-accuracy short-range as well, indicating a robust and versatile perception of the framework. Remarkably, the accuracy of OccFiner (VoxFormer) even surpasses onboard LiDAR-based SSC models including SSCNet, SSCNet-full, and LMSCNet. This is the first time that the camera-based SSC model surpasses the LiDAR-based SSC model in accuracy. Compared with the onboard LiDAR-based LMSCNet, the camera-based OccFiner (VoxFormer) achieves a relative mIoU improvement of $5.24\%$. This proves that OccFiner unleashes the potential of the camera-based SSC applications. This breakthrough in the camera-based versus LiDAR-based SSC model signifies a crucial advancement for autonomous driving, notably in the realms of pure visual SSC data closure and auto-labeling.
Table 3: Semantic scene completion results on the SemanticKITTI hidden test set [2]. * represents these methods are adapted for the RGB inputs, which are implemented and reported in MonoScene [6]. normal-$\dagger$ results from the official code release.
Method Modality
$\blacksquare$ car (3.92%)
$\blacksquare$ bicycle (0.03%)
$\blacksquare$ motorcycle (0.03%)
$\blacksquare$ truck (0.16%)
$\blacksquare$ other-vehicle (0.20%)
$\blacksquare$ person (0.07%)
$\blacksquare$ bicyclist (0.07%)
$\blacksquare$ motorcyclist (0.05%)
$\blacksquare$ road (15.30%)
$\blacksquare$ parking (1.12%)
$\blacksquare$ sidewalk (11.13%)
$\blacksquare$ other-ground (0.56%)
$\blacksquare$ building (14.1%)
$\blacksquare$ fence (3.90%)
$\blacksquare$ vegetation (39.3%)
$\blacksquare$ trunk (0.51%)
$\blacksquare$ terrain (9.17%)
$\blacksquare$ pole (0.29%)
$\blacksquare$ traffic-sign (0.08%)
IoU mIoU
LMSCNet* [52] Camera 14.30 0.00 0.00 0.30 0.00 0.00 0.00 0.00 46.70 13.50 19.50 3.10 10.30 5.40 10.80 0.00 10.40 0.00 0.00 31.38 7.07
3DSketch* [7] Camera 17.10 0.00 0.00 0.00 0.00 0.00 0.00 0.00 37.70 0.00 19.80 0.00 12.10 3.40 12.10 0.00 16.10 0.00 0.00 26.85 6.23
AICNet* [23] Camera 15.30 0.00 0.00 0.70 0.00 0.00 0.00 0.00 39.30 19.80 18.30 1.60 9.60 5.00 9.60 1.90 13.50 0.10 0.00 23.93 7.09
JS3C-Net* [70] Camera 20.10 0.00 0.00 0.80 4.10 0.00 0.20 0.20 47.30 19.90 21.70 2.80 12.70 8.70 14.20 3.10 12.40 1.90 0.30 34.00 8.97
OccFormer [82] Camera 21.60 1.50 1.70 1.20 3.20 2.20 1.10 0.20 55.90 31.50 30.30 6.50 15.70 11.90 16.80 3.90 21.30 3.80 3.70 34.53 12.32
MonoScene [6] Camera 18.80 0.50 0.70 3.30 4.40 1.00 1.40 0.40 54.70 24.80 27.10 5.70 14.40 11.10 14.90 2.40 19.50 3.30 2.10 34.16 11.08
TPVFormer [18] Camera 19.20 1.00 0.50 3.70 2.30 1.10 2.40 0.30 55.10 27.40 27.20 6.50 14.80 11.00 13.90 2.60 20.40 2.90 1.50 34.25 11.26
VoxFormer [27] Camera 21.80 0.90 1.20 5.60 2.80 1.00 1.30 0.00 54.10 24.70 27.80 9.20 23.40 14.20 24.60 9.00 23.40 6.70 6.60 43.09 13.60
OccFiner (Vox) Camera 31.10 1.90 1.60 7.80 2.50 3.80 2.00 0.00 62.00 35.20 36.10 13.70 28.10 18.70 29.10 9.60 31.70 6.60 6.50 45.54 17.27
w.r.t. VoxFormer - +9.30 +1.00 +0.40 +2.20 -0.30 +2.80 +0.70 - +7.90 +10.50 +8.30 +4.50 +4.70 +4.50 +4.50 +0.60 +8.30 -0.10 -0.10 +2.45 +3.67
SSCNet-Full [55] LiDAR 24.30 0.50 0.80 1.20 4.30 0.30 0.30 0.00 51.20 27.10 30.80 6.40 34.50 19.90 35.30 18.20 29.00 13.10 6.70 50.00 16.10
ESSCNet [80] LiDAR 26.40 0.30 5.40 5.00 9.10 2.90 2.70 0.10 43.80 26.90 28.10 10.30 29.80 23.30 35.80 20.10 28.70 16.40 16.70 41.80 17.50
LMSCNet [52] LiDAR 30.90 0.00 0.00 1.50 0.80 0.00 0.00 0.00 64.80 29.00 34.70 4.60 38.10 21.30 41.30 19.90 32.10 15.00 0.80 56.70 17.60
[…]
OccFiner (SCPNet) LiDAR 49.00 38.30 40.20 13.30 28.10 19.60 20.00 1.80 74.70 52.10 54.70 30.90 40.90 44.70 49.20 40.80 54.40 38.40 27.40 61.68 37.82
w.r.t. SCPNet - +6.70 +4.60 +6.60 +1.10 +2.10 +1.20 +3.00 +0.20 +5.20 +0.40 +5.20 +1.70 +5.50 +3.20 +7.40 +2.40 +4.70 +0.60 +0.30 +6.53 +3.28
As shown in Tab. 3, in the results submitted for the hidden test set of SemanticKITTI [2], OccFiner showcases remarkable performance. The OccFiner (Vox) and OccFiner (SCPNet) establish new state-of-the-art accuracy for camera-based and LiDAR-based SSC respectively. They deliver significant relative mIoU improvements compared to the best-published results of $26.99\%$ and $9.50\%$, respectively. The OccFiner (Vox) consistently surpasses the SSCNet-full [55]. This further demonstrates the efficacy and advancement of the OccFiner framework.
Table 4: Benchmarking results on SSCBench-KITTI360 [26]. The default evaluation range is $51.2{\times}51.2{\times}6.4m^{3}$.
Method
Input
IoU
mIoU
  car
  bicycle
  motorcycle
  truck
  other-veh.
  person
  road
  parking
  sidewalk
  other-grnd
  building
  fence
  vegetation
  terrain
  pole
  traf.-sign
  other-struct.
  other-object
LMSCNet [52] L 47.53 13.65 20.91 0.00 0.00 0.26 0.00 0.00 62.95 13.51 33.51 0.20 43.67 0.33 40.01 26.80 0.00 0.00 3.63 0.00
SSCNet [55] L 53.58 16.95 31.95 0.00 0.17 10.29 0.58 0.07 65.70 17.33 41.24 3.22 44.41 6.77 43.72 28.87 0.78 0.75 8.60 0.67
Voxformer [27] C 38.76 11.91 17.84 1.16 0.89 4.56 2.06 1.63 47.01 9.67 27.21 2.89 31.18 4.97 28.99 14.69 6.51 6.92 3.79 2.43
MonoScene [6] C 37.87 12.31 19.34 0.43 0.58 8.02 2.03 0.86 48.35 11.38 28.13 3.22 32.89 3.53 26.15 16.75 6.92 5.67 4.20 3.09
OccFiner (Mono) C 38.51 13.29 20.78 1.08 1.03 9.04 3.58 1.46 53.47 12.55 31.27 4.13 33.75 4.62 26.83 18.67 5.04 4.58 4.05 3.32
w.r.t. MonoScene - +0.64 +0.98 +1.44 +0.65 +0.45 +1.02 +1.55 +0.60 +5.12 +1.17 +3.14 +0.91 +0.86 +1.09 +0.68 +1.92 -1.88 -1.09 -0.15 +0.23
The effectiveness of OccFiner is also assessed on the SSCBench-KITTI360 dataset [26]. As shown in Tab. 4, OccFiner (Mono) demonstrates notable performance, achieving a relative improvement in mIoU of $7.96\%$ compared to MonoScene [6]. This underscores OccFiner’s capability to enhance SSC accuracy across diverse datasets.
4.3 Ablation Studies
Table 5: Ablation experiments. Settings used in our final framework are underlined. See Sec. 4.3 for details.
Experiment Method SemanticKITTI (val) #Parameters
Near Middle Far
Reference Model (VoxFormer [27]), Training: 10 Epoch (SemanticKITTI [2]).
Onboard Model VoxFormer [27] 22.13 18.58 13.33 551M
DualFlow4D Transformer w/o 21.71 19.23 13.96 +0.83M
with 23.29 20.25 14.52 +4.51M
Temporal Embedding Embedding 23.24 20.21 14.50 +4.90M
Not Embedding 23.29 20.25 14.52 +4.51M
Spatial Embedding Not Embedding 23.14 19.95 14.40 +4.12M
Embedding 23.29 20.25 14.52 +4.51M
Reference Frame w/o 22.99 20.04 14.30 +4.51M
with 23.29 20.25 14.52 +4.51M
Reference Model (VoxFormer [27]), Training: 10 Epoch (SemanticKITTI [2]) $\rightarrow$ Global Prop.
Offboard Baseline (Average Fusion) VoxFormer [27] + Vanilla Registration 21.18 18.36 14.52 551M
OccFiner Stage 1 + Vanilla Registration 22.60 20.19 16.11 +4.51M
Learning Refine w/o 21.48 18.97 16.05 n.a.
with 23.43 21.29 18.09 +4.51M
Frustum Weighting w/o 23.09 20.80 17.63 +4.51M
with 23.43 21.29 18.09 +4.51M
Distance-aware Weighting w/o 23.14 20.92 17.34 +4.51M
with 23.43 21.29 18.09 +4.51M
Temporal Input (-5, 5) 23.41 21.11 16.25 +4.51M
(-10, 10) 23.45 21.29 17.65 +4.51M
(-15, 15) 23.43 21.28 18.00 +4.51M
(-20, 20) 23.43 21.28 18.05 +4.51M
(-25, 25) 23.43 21.29 18.09 +4.51M
(-30, 30) 23.43 21.29 18.08 +4.51M
The efficacy of each module in the offboard OccFiner framework is quantified through ablation studies detailed in Tab. 5. Using VoxFormer [27] as the onboard component, we evaluate mIoU accuracy across different distances.
Multi-to-Multi Local Propagation Network. 1) DualFlow4D Transformer. We validate the critical role of the 4D Transformer for local SSC feature propagation, evidenced by a $0.56\%$ drop in mIoU upon its removal. 2) Temporal Embedding, however, did not positively impact accuracy. 3) Spatial Embedding. The result highlights the importance of spatial embedding and the introduction of a long-distance reference frame in the multi-to-multi paradigm, improving accuracy by $0.12\%$. 4) Reference Frame. OccFiner introduces a long-distance reference frame in addition to the local clip, which further improves the accuracy by $0.22\%$. It proves the positive effect of long-term geometric and semantic cues on error compensation. 5) Overall, OccFiner’s first stage, while adding few model parameters ($4.51$M vs. $551$M), effectively compensates for onboard model errors and propagates high-confidence features, boosting mIoU from $13.33\%$ to $14.52\%$, and uniformly improves mIoU across various distances. This is critical for the next global propagation stage.
Region-centric Global Propagation. 1) Average Fusion. Average fusion applied as a baseline to the onboard model yielded only a modest mIoU improvement at ‘far’, indicating limitations at near and middle distances with $0.95\%$ and $0.22\%$ mIoU drop, respectively. 2) Learning Refine. Removing the local propagation network resulted in a significant drop in mIoU by $2.04\%$, highlighting the critical role of first-stage error compensation in enhancing global propagation accuracy across various distances. 3) Sensor Bias. the frustum weighting and distance-aware weighting adjustments are shown to be vital, with their removal leading to noticeable accuracy drops. This underscores the importance of considering the camera field of view and depth uncertainties in long-distance measurements. 4) Temporal Input. The optimal number of frames for global propagation is identified around $(-25,25)$, with more frames offering no additional benefit. 5) Overall, learning-free global propagation significantly improves the accuracy of long-distance prediction. Compared with ‘OccFiner Stage 1 + Vanilla Registration’, OccFiner’s region-centric global propagation further improving the accuracy across different distances by $0.83\%$, $1.10\%$, and $1.98\%$.
5 Conclusions and Discussion
In addressing the challenges of inferior performance and data closure in vision-based SSC, we introduce OccFiner, the first offboard SSC setup to solve the unreliability of the onboard model. By removing computational constraints, OccFiner can reason all frames together and build multi-view consistent SSC predictions. Concretely, OccFiner leverages a hybrid propagation strategy. In the multi-to-multi local propagation stage, OccFiner focuses on the onboard model error compensation and implicit semantic and geometric cues propagation. In the region-centric global propagation stage, OccFiner considers sensor bias and manages long-term information aggregation. Our experimental findings validate OccFiner’s versatility across various onboard models and its capability to significantly elevate SSC quality, setting new benchmarks in vision- and LiDAR-based algorithms on SemanticKITTI. We hope that our framework can become a potent augmentor for onboard SSC algorithms and inspire future offboard research.
Limitations. Our offboard framework, OccFiner, achieves notable results but requires individual retraining for each onboard model. This process is tailored to address the specific errors of each onboard model, with its effectiveness partly dependent on the scene’s characteristics. Additionally, there is substantial scope for improving offboard SSC accuracy in comparison to the ground truth.
Broader impact, Ethics. While OccFiner improves 3D scene understanding, potential inaccuracies in its predictions could have serious implications, especially in critical applications like autonomous driving. It is crucial that such technologies be complemented by additional safety measures to mitigate risks associated with these inevitable errors.